# Batch learning equals online learning in Bayesian supervised learning

### v1

@@ 3. Bayesian inversions in Bayesian supervised learning models
Completion of the proof of Theorem 3.4. The first and second assertion of Theorem 3.4 follow immediately from Lemma 3.5. ∎
@@
Here $Dir_{(A)}(\beta)$ is the Dirichlet distribution with parameter $\beta\in{\mathcal{M}}^{*}({\Omega}_{(A)})$, and $M_{*}(\pi_{(A)}):{\mathcal{M}}^{*}({\mathcal{Y}})\to{\mathcal{M}}({\Omega}_{{}_{|(A)|}})$ is defined by the same formula (2.2). Let us consider a Bayesian statistical model $\big({\mathcal{P}}({\mathcal{Y}}),{\mathcal{D}}(\alpha),\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})},{\mathcal{Y}}\big)$ associated with the case #(X)=1\#(X)=1 and $k=1$ in Theorem 3.4. Lemma 3.5 and the above commutative diagram imply that the condition of Theorem 3.4 holds for $\big({\mathcal{P}}({\mathcal{Y}}),{\mathcal{D}}(\alpha),\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})},{\mathcal{P}}({\mathcal{Y}})\big)$. Using Bayes’ formula, one computes that
@@
For $(A)\leq(B)\in(\pi({\mathcal{Y}}),\leq)$, we can immediately verify that the following diagram
@@
$\mathrm{\operatorname{Id}}^{n}_{{\mathcal{P}}({\mathcal{X}})}:{\mathcal{P}}({\mathcal{Y}})\to{\mathcal{P}}({\mathcal{Y}}^{n}),\mu\mapsto\otimes^{n}\mu$
@@ 4. Posterior predictive distributions
Let $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ be a Bayesian model for supervised learning, $S_{n}=(x_{1},\ldots,x_{n})\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, and $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$.
1) Let ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ be a regular conditional probability measure for the joint distribution
$\mu^{0}_{T_{m},S_{n},\mu_{\Theta}}:=\underline{\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{\Pi_{\mathcal{X}}(S_{n})})}_{*}\mu_{\Theta}\in{\mathcal{P}}({\mathcal{Y}}^{m}\times{\mathcal{Y}}^{n}).$
Then ${\mathbf{q}}^{n}_{m}\big(\Pi_{\mathcal{Y}}(S_{n})\big)$ is the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ of the tuple $(y_{1}^{\prime},\ldots y^{\prime}_{m})$ where $y^{\prime}_{i}$ is the label of $t_{i}$, $i\in\overline{1,m},$ after seeing $S_{n}$  [^1] See Remark 2.2 for the uniqueness of ${\mathbf{q}}^{n}_{m}$..
2) For $1\leq k\leq n$ we let $S_{k}:=\big((x_{1},y_{1}),\ldots,(x_{k},y_{k})\big)$. The predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ can be computed recursively as follows.
@@
    Step 1: Let ${\mathbf{q}}^{1}_{m+n-1}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}^{m+n-1})$ be a regular conditional probability measure for the joint distribution $\mu^{0}_{T_{m},S_{n},\mu_{\Theta})}\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$ with respect to the projection ${\mathcal{Y}}^{m+n}\to{\mathcal{Y}}$. Then we set
    (4.1) $\mu_{(T_{m},S_{n},\mu_{\Theta})}^{1}:={\mathbf{q}}^{1}_{m+n-1}(y_{n})\in{\mathcal{P}}({\mathcal{Y}}^{m+n-1}).$
@@
    Step $k+1$ for $1\leq k\leq n-1$. Let $Pr^{m+n-k}_{m+n-k-1}:{\mathcal{Y}}^{m+n-k}\to{\mathcal{Y}}^{m+n-k-1}$ be the projection on the first (m+nk1)(m+n-k-1) factors. Let ${\mathbf{q}}^{k+1}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}^{m+n-k-1})$ be a regular conditional probability measure for the joint distribution $\mu^{k}_{T_{m},S_{n},\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k})$. Then we set
    (4.2) $\mu_{(T_{m},S_{n},\mu_{\Theta})}^{k+1}:={\mathbf{q}}^{k}(y_{n-k})\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k-1}).$
Then $\mu_{T_{m},S_{n},\mu_{\Theta}}^{n}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ is the posterior predictive distribution of ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}$.
@@
2) To prove the second assertion of Theorem 4.1, we consider the following diagram
@@
$\mu^{n-1}_{(T_{m},S_{n},\mu_{\Theta})}={\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}.$
For $X_{m}:=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$ we denote by $X_{m}]$ the smallest subset of $\mathcal{X}$} that contains each of $x_{i}$. From Theorem 4.1 we obtain immediately the following.
@@
Let $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ be a universal Bayesian supervised models. Let $T_{m}\in{\mathcal{X}}^{m},S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$ and $A:=[T_{m}]\cup[\Pi_{\mathcal{X}}(S_{n})]$. Let $R_{A}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{A},h\mapsto h_{|A},$ denote the natural restriction map. Then we have
(4.3) ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}={\mathcal{P}}_{T_{m}|S_{n},(R_{A})_{*}(\mu_{\Theta})}.$
@@
as a universal Bayesian supervised learning model for learning the corrupted measurement (4.4).
@@
be the Markov kernel describing the sampling distribution of uncorrupted measurement. By Proposition 3.3, for $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$, the predictive distribution of the tuple $(f(t_{1}),\ldots,f(t_{m}))$ after seeing $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$ can be chosen as the value ${\mathbf{q}}^{n}_{m}(\Pi_{\mathcal{Y}}(S_{n}))\in{\mathcal{P}}(V^{m})$ where ${\mathbf{q}}^{n}_{m}:V^{m}\to V^{n}$ is a regular conditional probability measure for the joint distribution $(\underline{\mathfrak{m}^{2}(\mathfrak{m}^{m}\circ E_{T_{m}}\circ{\mathbf{p}}^{0}\circ h,\mathfrak{m}^{n}\circ\Pi_{\Pi_{\mathcal{X}}(S_{n})}\circ{\mathbf{p}}^{\varepsilon}\circ h)})_{*}\mu_{\Theta}\in{\mathcal{P}}(V^{m}\times V^{n})$.
@@
Theorem 4.3 (Posterior predictive distribution with corrupted measurement).
Let $(\Theta,\mu_{\Theta},h,V^{\mathcal{X}})$ be a Bayesian model for regression learning, $S_{n}=\big((x_{1},y_{1})\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times V)^{n}$ be training data with $y_{i}$ being a corrupted measurement of $f(x_{i})$ for $i\in\overline{1,n}$, and $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$.
1) Let ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ be a regular conditional probability measure for the joint distribution
$\mu^{0}_{(T_{m},S_{n},\mu_{\Theta})}:=\underline{\mathfrak{m}^{2}(h^{0}_{T_{m}},h^{\varepsilon}_{\Pi_{\mathcal{X}}(S_{n})})}_{*}\mu_{\Theta}\in{\mathcal{P}}(V^{m}\times V^{n}).$
Then ${\mathbf{q}}^{n}_{m}\big(\Pi_{V}(S_{n})\big)$ is the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$ of the tuple $(y_{1}^{\prime},\ldots y^{\prime}_{m})$ where $y^{\prime}_{i}\in V$ is the label of $t_{i}$, $i\in\overline{1,m},$ after seeing $S_{n}=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$.
2) For $1\leq k\leq n$ we let $S_{k}:=\big((x_{1},y_{1}),\ldots,(x_{k},y_{k})\big)$. The posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$ can be computed recursively as follows.
@@
    Step 1: Let ${\mathbf{q}}^{1}:V\to{\mathcal{P}}(V^{m+n-1})$ be a regular conditional probability measure for the joint distribution $\mu^{0}_{(T_{m},S_{n},\mu_{\Theta})}\in{\mathcal{P}}(V^{m+n})$ with respect to the projection $V^{m+n}\to V$. Then we set
    (4.6) $\mu_{(T_{m},S_{n},\mu_{\Theta})}^{1}:={\mathbf{q}}^{1}(y_{n})\in{\mathcal{P}}(V^{m+n-1}).$
@@
    Step $k+1$ for $1\leq k\leq n-1$. Let $Pr^{m+n-k}_{m+n-k-1}:V^{m+n-k}\to V^{m+n-k-1}$ be the projection on the first (m+nk1)(m+n-k-1) factors. Let ${\mathbf{q}}^{k+1}:V\to{\mathcal{P}}(V^{m+n-k-1})$ be a regular conditional probability measure for the joint distribution $\mu^{k}_{(T_{m},S_{n},\mu_{\Theta})}\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k})$. Then we set
    (4.7) $\mu_{(T_{m},S_{n},\mu_{\Theta})}^{k+1}:={\mathbf{q}}^{k}(y_{n-k})\in{\mathcal{P}}(V^{m+n-k-1}).$
Then $\mu_{(T_{m},S_{n},\mu_{\Theta})}^{n}\in{\mathcal{P}}(V^{m})$ is the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$.
@@
We illustrate Theorem 4.3 with Gaussian process regression model $({\mathbb{R}}^{\mathcal{X}},{\mathcal{G}}{\mathcal{P}}(m,K),{\mathbf{p}}^{\varepsilon},{\mathcal{P}}({\mathbb{R}})^{\mathcal{X}})$, where ${\mathcal{G}}{\mathcal{P}}(m,k)$ is a Gaussian measure on $V^{\mathcal{X}}$ defined by the mean function $m\in{\mathbb{R}}^{\mathcal{X}}$ and $K:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$ is a positive definite kernel. Let $\mu$ be a Gaussian measure on the function space $V^{\mathcal{X}}$ where $\mathcal{X}$} is an input space and $V={\mathbb{R}}^{k}$. One sees immediately that the recipe for computing the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{G}}{\mathcal{P}}(m,K)}$ in Theorem 4.3(1) coincides with the classical formula for posterior predictive distributions in Gaussian process regression described in [RW2006]. Furthermore, the recursive formula in Theorem 4.3(2) is much simpler and faster than the classical formula since it does not require computing the (pseudo) inverse of a square matrix of size $(n\times n)$ associated with the kernel $K^{\varepsilon}_{n}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ which is the variance of the Gaussian measure that governs the distribution of $y_{i}=f(x_{i})+{\varepsilon}(x_{i})\in{\mathbb{R}}$, $i\in\overline{1,n}$, and a multiplication of matrices of size $(m\times n)$ with this pseudo inverse and with a matrix of size $(n\times m)$ [Stein1999, Appendix A]. Instead, we have to compute $n$-round of multiplications of matrices of size $m+n-i\times 1$ with a matrix of size $1\times m+n-i$ for $i\in\overline{1,k}$. This sequential update procedure is known to be equivalent to the celebrated Kalman filter update equations, see, e.g., [SS2023, §6.3].

### v2

@@ 3. Bayesian inversions in Bayesian supervised learning models
Completion of the proof of Theorem 3.4. The measurability of ${\mathbf{q}}_{X_{k}}$ follows from Lemma 3.5, taking into account the commutative diagram 3.14. To complete the proof of Theorem 3.4, we need to prove (3.15). By Lemma 3.5, the RHS of (3.15) equals $(\Gamma_{\underline{\mathfrak{m}^{k}\circ E_{X_{k}}}}P_{*}(\pi_{(\tilde{A})})^{\mathcal{X}}_{*}\mu({\mathcal{S}}_{1}\times\ldots\times{\mathcal{S}}_{n}\times B)$ and the LHS of (3.15) equals $(\Gamma_{\underline{{\mathbf{q}}^{(\tilde{A})}_{X_{k}}}})_{*}(\underline{\mathfrak{m}^{k}\circ E_{X_{k}}})_{*}P_{*}(\pi_{(\tilde{A})})^{\mathcal{X}}_{*}\mu$. Since ${\mathbf{q}}^{(\tilde{A})}_{X_{k}}$ is a Bayesian inversion of $\mathfrak{m}^{k}\circ E_{X_{k}}$, (3.15) holds. ∎
@@
Here $Dir_{(A)}(\beta)$ is the Dirichlet distribution with parameter $\beta\in{\mathcal{M}}^{*}({\Omega}_{(A)})$, and $M_{*}(\pi_{(A)}):{\mathcal{M}}^{*}({\mathcal{Y}})\to{\mathcal{M}}({\Omega}_{{}_{|(A)|}})$ is defined by the same formula (2.2). Let us consider a Bayesian statistical model $\big({\mathcal{P}}({\mathcal{Y}}),{\mathcal{D}}(\alpha),\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})},{\mathcal{Y}}\big)$ associated with the case #(X)=1\#(X)=1 and $k=1$ in Theorem 3.4. Lemma 3.5 and the above commutative diagram imply that the condition of Theorem 3.4 holds for $\big({\mathcal{P}}({\mathcal{Y}}),{\mathcal{D}}(\alpha),\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})},{\mathcal{P}}({\mathcal{Y}})\big)$. It is known that
@@
For $(A)\leq(B)\in(\pi({\mathcal{Y}}),\leq)$, we verify immediately that the following diagram
@@
$\mathrm{\operatorname{Id}}^{n}_{{\mathcal{P}}({\mathcal{Y}})}:{\mathcal{P}}({\mathcal{Y}})\to{\mathcal{P}}({\mathcal{Y}}^{n}),\mu\mapsto\otimes^{n}\mu$
@@ 4. Posterior predictive distributions
Let $(\Theta,\mu_{\Theta},{\mathbf{p}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ be a Bayesian model for supervised learning, $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$, and $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$. For $Y_{n}=(y_{1},\ldots,y_{n})\in{\mathcal{Y}}^{n}$ let $S_{n}(X_{n},Y_{n}):=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$.
1) The posterior predictive distribution ${\mathcal{P}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ defined by Equation (2.10) is a regular conditional probability measure for the joint distribution
$\mu^{0}_{T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta}}:=\underline{\mathfrak{m}^{2}({\mathbf{p}}_{T_{m}},{\mathbf{p}}_{X_{n}})}_{*}\mu_{\Theta}\in{\mathcal{P}}({\mathcal{Y}}^{m}\times{\mathcal{Y}}^{n})$
with respect to the projection $\Pi_{{\mathcal{Y}}^{n}}:{\mathcal{Y}}^{m}\times{\mathcal{Y}}^{n}\to{\mathcal{Y}}^{n}$. Hence, if ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ is a regular conditional probability measure for $\mu^{0}_{T_{m},S_{n}(X_{m}Y_{n}),\mu_{\Theta}}$ with respect to the projection $\Pi_{{\mathcal{Y}}^{n}}$ then ${\mathbf{q}}^{n}_{m}\big(Y_{n}\big)$ coincides with ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ up to a $(\underline{{\mathbf{p}}_{X_{n}}})_{*}\mu_{\Theta}$-zero set.
2) The posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ can be computed recursively as follows.
@@
    Step 1: Let ${\mathbf{q}}^{1}_{m+n-1}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}^{m+n-1})$ be a regular conditional probability measure for the joint distribution $\mu^{0}_{T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$ with respect to the projection ${\mathcal{Y}}^{m+n}\to{\mathcal{Y}}$. Then we set
    (4.1) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{1}:={\mathbf{q}}^{1}_{m+n-1}(y_{n})\in{\mathcal{P}}({\mathcal{Y}}^{m+n-1}).$
@@
    Step $k+1$ for $1\leq k\leq n-1$. Let ${\mathbf{q}}^{k+1}:{\mathcal{Y}}\to{\mathcal{P}}({\mathcal{Y}}^{m+n-k-1})$ be a regular conditional probability measure for the joint distribution $\mu^{k}_{T_{m},S_{n},\mu_{\Theta}}\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k})$ with respect to the projection ${\mathcal{Y}}^{m+n-k}\to{\mathcal{Y}}$. Then we set
    (4.2) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{k+1}:={\mathbf{q}}^{k}(y_{n-k})\in{\mathcal{P}}({\mathcal{Y}}^{m+n-k-1}).$
Then $\mu_{T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta}}^{n}\in{\mathcal{P}}({\mathcal{Y}}^{m})$ is the posterior predictive distribution of ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}$.
@@
2) For $1\leq k\leq n$ we let $S_{k}:=\big((x_{1},y_{1}),\ldots,(x_{k},y_{k})\big)$. To prove the second assertion of Theorem 4.1, we consider the following diagram
@@
$\mu^{n}_{(T_{m},S_{n},\mu_{\Theta})}={\mathcal{P}}_{T_{m}|S_{n},\mu_{\Theta}}.$
For $X_{m}:=(x_{1},\ldots,x_{m})\in{\mathcal{X}}^{m}$ we denote by $X_{m}]$ the smallest subset of $\mathcal{X}$} that contains each of $x_{i}$.
@@
Let $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ be a universal Bayesian supervised models. Let $T_{m}\in{\mathcal{X}}^{m},X_{n}\in{\mathcal{X}}^{n}$ and $A:=[T_{m}]\cup[X_{n}]$. Let $R_{A}^{\mathcal{X}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{A},h\mapsto h_{|A},$ denote the natural restriction map. Denote by $\operatorname{Id}_{X_{n}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{n})$ the composition $\mathfrak{m}^{n}\circ E_{X_{n}}$. Then for (weakly) $(\underline{\operatorname{Id}_{X_{n}}})_{*}\mu$-a.s. $Y_{n}\in{\mathcal{Y}}^{n}$ we have
(4.3) ${\mathcal{P}}_{T_{m}|S_{n}(X_{,}Y_{n}),\mu}={\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),(R_{A}^{\mathcal{X}})_{*}\mu},$
where the RHS of (4.3) is the posterior predictive distribution of the restricted Bayesian supervised learning model $({\mathcal{P}}({\mathcal{Y}})^{A},(R_{A}^{\mathcal{X}})_{*}\mu_{\Theta},\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{A}},{\mathcal{P}}({\mathcal{Y}})^{A})$.
Proof.
We consider the following diagram
[…]
where $\operatorname{Id}_{T_{m}}:=\mathfrak{m}^{m}\circ E_{T_{m}}:{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ and $\operatorname{Id}^{A}_{T_{m}}:=\mathfrak{m}^{m}\circ E_{T_{m}}:{\mathcal{P}}({\mathcal{Y}})^{A}\to{\mathcal{P}}({\mathcal{Y}}^{m})$. Similarly, we define $\operatorname{Id}^{A}_{X_{n}}$. By Theorem 4.1, ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},\cdot),\mu}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ is a regular conditional probability measure for the joint distribution $(\underline{\operatorname{Id}_{T_{m},X_{n}}})_{*}\mu\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$, and ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},\cdot),(R_{A}^{\mathcal{X}})_{*}\mu}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ is a regular conditional probability measure for the joint distribution $(\underline{\operatorname{Id}^{A}_{T_{m},X_{n}}})_{*}(R_{A})_{\mu}\in{\mathcal{P}}({\mathcal{Y}}^{m+n})$. To conclude Corollary 4.2, we note that
$(\underline{\operatorname{Id}_{T_{m},X_{n}}})_{*}\mu=(\underline{\operatorname{Id}^{A}_{T_{m},X_{n}}})_{*}(R_{A}^{\mathcal{X}})_{\mu}$
since $T_{m}]\subset A$ and $X_{n}]\subset A$, hence $\operatorname{Id}_{T_{m},X_{n}}=\operatorname{Id}^{A}_{T_{m},X_{n}}\circ R_{A}^{\mathcal{X}}$. Therefore, both the Bayesian statistical supervised model $({\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}},\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}})$ and the Bayesian statistical supervised model $({\mathcal{P}}({\mathcal{Y}}^{A},(R^{\mathcal{X}}_{A})_{*}\mu,\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{A}})$ induce the same joint distribution on ${\mathcal{P}}({\mathcal{Y}}^{m+n})$. Taking into account Theorem 4.1, we obtain Corollary 4.2. ∎
@@
as a universal Bayesian supervised learning model for learning the corrupted measurements (4.5).
@@
be the Markov kernel describing the sampling distribution of uncorrupted measurement. By Proposition 3.3, for $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$, the predictive distribution of the tuple $(f(t_{1}),\ldots,f(t_{m}))$ after seeing $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$ can be chosen as the value ${\mathbf{q}}^{n}_{m}(\Pi_{\mathcal{Y}}(S_{n}))\in{\mathcal{P}}(V^{m})$ where ${\mathbf{q}}^{n}_{m}:V^{m}\to V^{n}$ is a regular conditional probability measure for the joint distribution $(\underline{\mathfrak{m}^{2}(\mathfrak{m}^{m}\circ E_{T_{m}}\circ{\mathbf{p}}^{0}\circ h,\mathfrak{m}^{n}\circ E_{\Pi_{\mathcal{X}}(S_{n})}\circ{\mathbf{p}}^{\varepsilon}\circ h)})_{*}\mu_{\Theta}\in{\mathcal{P}}(V^{m}\times V^{n})$.
@@
Theorem 4.3 (Posterior predictive distribution with corrupted measurements).
Let $(\Theta,\mu_{\Theta},h,V^{\mathcal{X}})$ be a Bayesian model for regression learning, with $y_{i}$ being a corrupted measurement of $f(x_{i})$ for $i\in\overline{1,n}$, and $T_{m}=(t_{1},\ldots,t_{m})\in{\mathcal{X}}^{m}$. For $X_{n}=(x_{1},\ldots,x_{n})\in{\mathcal{X}}^{n}$ and $Y_{n}=(y_{1},\ldots,y_{n})\in V^{n}$ let $S_{n}(X_{n},Y_{n}):=\big((x_{1},y_{1}),\ldots,(x_{n},y_{n})\big)\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$.
1) The posterior predictive distribution
${\mathcal{P}}^{n}_{m}:V^{n}\to{\mathcal{P}}(V_{m}),\,Y_{n}\mapsto{\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}},$
is a regular conditional probability measure for the joint distribution
$\mu^{0}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}:=\underline{\mathfrak{m}^{2}(h^{0}_{T_{m}},h^{\varepsilon}_{\Pi_{\mathcal{X}}(S_{n})})}_{*}\mu_{\Theta}\in{\mathcal{P}}(V^{m}\times V^{n})$
with respect to the projection $\Pi_{V^{n}}:V^{m}\times V^{n}\to V^{n}$, Hence, if ${\mathbf{q}}^{n}_{m}:V^{n}\to{\mathcal{P}}(V^{m})$ is a regular conditional probability measure for the joint distribution $\mu^{0}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}$ with respect to the projection $\Pi_{V^{n}}$ then ${\mathbf{q}}^{n}_{m}\big(\Pi_{V}(S_{n})\big)$ coincides with ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$ up to a $(\underline{h^{\varepsilon}_{X_{n}}})_{*}\mu_{\Theta}$-zero set.
2) The posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$ can be computed recursively as follows.
@@
    Step 1: Let ${\mathbf{q}}^{1}:V\to{\mathcal{P}}(V^{m+n-1})$ be a regular conditional probability measure for the joint distribution $\mu^{0}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}\in{\mathcal{P}}(V^{m+n})$ with respect to the projection $V^{m+n}\to V$. Then we set
    (4.7) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{1}:={\mathbf{q}}^{1}(y_{n})\in{\mathcal{P}}(V^{m+n-1}).$
@@
    Step $k+1$ for $1\leq k\leq n-1$. Let ${\mathbf{q}}^{k+1}:V\to{\mathcal{P}}(V^{m+n-k-1})$ be a regular conditional probability measure for the joint distribution $\mu^{k}_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}\in{\mathcal{P}}(V^{m+n-k})$ with respect to the projection $V^{m+n-k}\to V$. Then we set
    (4.8) $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{k+1}:={\mathbf{q}}^{k}(y_{n-k})\in{\mathcal{P}}(V^{m+n-k-1}).$
Then $\mu_{(T_{m},S_{n}(X_{n},Y_{n}),\mu_{\Theta})}^{n}\in{\mathcal{P}}(V^{m})$ is the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n}(X_{n},Y_{n}),\mu_{\Theta}}\in{\mathcal{P}}(V^{m})$.
@@
We illustrate Theorem 4.3 with Gaussian process regression model $({\mathbb{R}}^{\mathcal{X}},{\mathcal{G}}{\mathcal{P}}(m,K),{\mathbf{p}}^{\varepsilon},{\mathcal{P}}({\mathbb{R}})^{\mathcal{X}})$, where ${\mathcal{G}}{\mathcal{P}}(m,k)$ is a Gaussian measure on $R^{\mathcal{X}}$ defined by its mean function $m\in{\mathbb{R}}^{\mathcal{X}}$ and its covariance function $K:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$, which is a positive definite kernel. Let $\mu$ be a Gaussian measure on the function space $R^{\mathcal{X}}$ where $\mathcal{X}$} is an input space and $V={\mathbb{R}}$. One sees immediately that the recipe for computing the posterior predictive distribution ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{G}}{\mathcal{P}}(m,K)}$ in Theorem 4.3(1) coincides with the classical formula for posterior predictive distributions in Gaussian process regression described in [RW2006]. Furthermore, the recursive formula in Theorem 4.3(2) is much simpler and faster than the classical formula since it does not require computing the (pseudo) inverse of a square matrix of size $(n\times n)$ associated with the kernel $K^{\varepsilon}_{n}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ which is the variance of the Gaussian measure that governs the distribution of $y_{i}=f(x_{i})+{\varepsilon}(x_{i})\in{\mathbb{R}}$, $i\in\overline{1,n}$, and a multiplication of matrices of size $(m\times n)$ with this pseudo inverse and with a matrix of size $(n\times m)$ [Stein1999, Appendix A], [Rao2002, Chapter 8]. Instead, we have to compute $n$-round of multiplications of matrices of size $m+n-i\times 1$ with a matrix of size $1\times m+n-i$ for $i\in\overline{1,n}$. This sequential update procedure is known to be equivalent to the celebrated Kalman filter update equations, see, e.g., [SS2023, §6.3].